# The hierarchies of identities and closed products for multiple complexes

### v2

@@ 2. Multiple complexes and the main result
Indeed, each a differential identity $(\chi)=0$ from the system (2.8) can be “integrated“ with respect the differentials $d$ or $\overline{d}$ to find a multiple product $(\Gamma)$ such that $d_{a}(\Gamma)=(\chi)=0$. When we say integrated we mean that for an element $\gamma\in\mathcal{C}$ we find an element $\gamma^{\prime}$ such that $\gamma^{\prime}=d_{a}\gamma$. In addition to that relations specified by the identities (2.8) can be used in order to derive closed products which do not follow explicitly from the integration of identities. Let us underline that we work with a multiple complex spaces as grading subspaces of an associative algebra. In particular, the multiple product $\cdot_{l}$ can be introduces for geometric objects of general kinds. Thus, we keep the constructions of this paper independent of commutation relations. In the next three Sections we will provide instructive examples of identities and closed products of the first order of differentials actions. The full description and classification of differential identities and closed products at higher orders and powers of differentials action will be given in a forthcoming paper. The identities and closed products given in Sections (3)–5 generalize to symmetrized versions with respect to permutations of indices, positions of $\mathcal{C}$-elements, and orders and powers of differentials.
@@ 3. Examples of identities for multiple products
In this Section we illustrate the proof of Theorem 1 by complexity-growing examples. In this and next two Sections we always assume that all completions are closed with respect to corresponding differentials.
@@ 3.1. Two differentials, and the total maximal power two for two kinds of elements
00 0 $=$ $d_{a}\left(\ldots,(\phi)_{i},\ldots,(\psi)_{j},\ldots\right)$
@@ 3.3. Two differentials, identical elements of the total maximal power
For $q(\phi)$ identical elements $\phi$, placed at all $i$-th positions, $1\leq i\leq q(\phi)\in\mathbb{N}$, such that $\phi\in\mathcal{I}(q(\phi))$, i.e., $\phi^{q(\phi)}=0$, equivalently, (($\ldots$( …, $(\phi)_{1}$, $\ldots$, $(\phi)_{i}$, $\ldots$, $(\phi)_{q(\phi)},\ldots)=0$, (i.e., $\phi$ is placed $q(\phi)$ times in various places in the product). Then,
@@ 3.4. Orders of two differentials of several identical elements
$=\sum\limits_{i=1}^{k}\sum\limits_{s_{1}=1}^{q(p_{i},\phi)}\left(\ldots,\left(d_{a}^{p_{i}+1}\phi\right)_{s_{1}},\ldots,\left(d_{a}^{p_{i}}\phi\right)^{q(p_{i})-1}_{\widehat{s_{1}}},\ldots,\left(d_{a}^{p_{\widehat{i}}}\phi\right)^{q_{\widehat{i}}},\ldots\right),$
@@ 3.5. Two differentials, multiple elements
Consider the case of a single differential and multiple elements. In order to form the vanishing product, we include exactly one maximal power $q(\phi_{i})$ of an element $\phi_{i}$, and powers $1\leq r_{\widehat{i}}<q(\phi_{\widehat{i}})$ of other types of elements $\phi_{\widehat{i}}$, $11\leq{\widehat{i}}\neq i\leq k-$ 1. For $d_{a}$ as before
00 0 $=$ $d_{a}\sum\limits_{i=1}^{k}\left(\ldots,\phi_{1}^{r_{1}},\ldots,\phi_{i}^{q(\phi_{i})},\ldots,(\phi_{k})^{r_{k}},\ldots\right)$ (3.6)
$=$ $\sum\limits_{i=1}^{k}\sum\limits_{s=1}^{q(\phi_{i})}\left(\ldots,(d_{a}\phi_{i})_{s},\ldots,(\phi_{i})^{q(\phi_{i})-1}_{\widehat{s}},\ldots,\phi_{\widehat{i}}^{q(\phi_{\widehat{i}})},\ldots\right).$
@@
$=\sum\limits_{i=1}^{k}\sum\limits_{s_{1}=1}^{q(p_{i},\phi_{i})}\left(\ldots,(d_{a}^{p_{i}(\phi_{i})+1}\phi_{i})_{s_{1}},\ldots,(d_{a}^{p_{i}}\phi_{i})^{q_{i}(p_{i},\phi_{i})-1}_{\widehat{s_{1}}},\ldots,(d_{a}^{p_{\widehat{i}}}\phi_{\widehat{i}})^{r_{\widehat{i}}},\ldots\right).$
@@
One can also take the case of the total maximal power elements. I.e., $\sum_{i=1}^{t}$ $q_{i}$ $(p_{i},\phi_{i})=q$, and $q_{i}(p_{i},\phi_{i})<q(p_{i},\phi_{i})$ i.e., when the sum of powers reaches the critical value, and ×$\times\left(d_{a}^{p_{i}(\phi_{i})}\phi\right)^{q_{i}(\phi_{i})}\in\mathcal{I}(q)$. Similarly, $\sum_{i=1}^{t}p_{i}(\phi_{i})=p$, and $p_{i}(\phi_{i})<p(\phi_{i})$, i.e., the sum of orders of the differentials is at its maximum, and ×$\times d_{a}^{p_{i}(\phi_{i})}\phi_{i}\in\mathcal{I}(p)$ Then in both cases $\left(\ldots,\left(d_{a}^{p_{i}(\phi_{i})}\phi_{i}\right)^{q_{i}(p_{i},\phi_{i})},\ldots\right)=0$. When we apply the second differential $d_{\overline{a}}$ to the identity (3.7), all branches of the hierarchy of identities arise with appropriate conditions on orders and powers of differentials.
@@ 4. Examples of a single element closed products
In this Section we provide examples of closed products arising, in particular, from the identities of Section 3. Closed products and their identities are very useful in the theory of completely integrable [2] and exactly solvable [16, 18] systems. Note also possible applications for operads. Note that in the search of products annihilated by a differential, we take only non-vanishing products. To use the annihilation given by maximal powers of $d_{a}\phi_{i}$ in multi-element products after application of a differential $d_{a}$ once, we have to include not just one power of each element but lower powers also. The main idea of a closed product construction is in taking into account changes both in orders of differentials themselves and powers of their action on $\mathcal{C}$-elements. For that reason we have to include not only powers of elements acted by the differentials, but also powers of elements not acted by them. The balance of raising and lowering powers of elements results in extra conditions producing closed products. In some of the examples below we use the transfer of differentials using certain identities from Section 3. By integrating the identities of the form (3.5) one finds sets of advanced closed products.
@@ 4.3. Orders of a single differential, single element, powers of differentials
$d_{a}\left(\ldots,\left(d_{a}^{p_{i}}\phi\right)^{q_{i}}_{i},\ldots\right)$ $=$ $\sum\limits_{s=1}^{k}\sum\limits_{t=1}^{q_{s}}\left(\ldots,\left(d_{a}^{p_{s}+1}\phi\right)_{t},\ldots,\left(d_{a}^{p_{s}}\phi\right)^{q_{s}-1}_{\widehat{t}},\ldots,\left(d_{a}^{p_{\widehat{s}}}\phi\right)^{q_{\widehat{s}}}_{\widehat{s}},\ldots\right),$
@@ 5.4. Powers of orders of multiple differentials, multiple elements
when 1.) there exists $i$, $1\leq i\leq k$, such that $p_{n+1,i}\left(d^{p_{n,i}}_{a_{n,i}}\right.$ldots $d^{p_{1},i}_{a_{1,i}}$ $\left.\phi\right)$geq $p\left(d^{p_{n+1,i}}_{a_{n,i}}\ldots d^{p_{1},i}_{a_{1,i}}\phi\right)$; 2.) if $k\geq 1$, and $({\bf p}_{n+1,i},{\bf a}_{n+1,i})=({\bf p}_{n,j},{\bf a}_{n,j})$, for some $1\leq i\neq j\leq k$.
@@ 6. Multiple-graded differential algebras
where it is clear that $q_{s,i}<q\left(\mathcal{D}_{\bm{J}_{s,i}}\phi_{s,i}\right)$ are less than the maximal powers of the corresponding elements $\mathcal{D}_{\bm{J}_{s,i}}\phi_{s,i}$, and the maximal orders of differentials constituting (as in (2.6)) $\mathcal{D}_{\bm{J}_{s,i}}\phi_{s,i}$ satisfy the recurrence conditions (2.7). We call this a differential condition. Symmetrizing with respect to all choices of ${\bm{J}_{s}}$ for a fixed $l$, we obtain a set of differential conditions
@@
for $\gamma_{s}\in\mathcal{C}$, $s\in\mathbb{Z}$. With $\mathcal{D}_{\bm{J}_{s}}\gamma_{s}=0$ for some $s$, we call (6.2) an orthogonality condition. For each differential condition above of the set (6.2) we have the coherence condition for corresponding upper and lower indices of $\mathcal{C}$-spaces, i.e., ${\bm{n}}$ $\left(\mathcal{D}_{J_{s}}\gamma_{s}\right)$ $=$ $\sum_{i=1}^{l}\left({\bm{n}_{i}}\left(\mathcal{D}_{\bm{J}_{s,i}}\phi_{s,i}\right)-r_{i}\left(\mathcal{D}_{\bm{J}_{s,i}}\phi_{s,i}\right)\right)$, ${\bm{m}}\left(\mathcal{D}_{\bm{J}_{s}}\gamma_{s}\right)=\sum_{i=1}^{l}\left({\bm{m}}_{i}\left(\mathcal{D}_{\bm{J}_{s,i}}\phi_{s,i}\right)\right.$ $\left.-{\bm{t}}_{i}\left(\mathcal{D}_{\bm{J}_{s,i}}\phi_{s,i}\right)\right)$. We will skip these coherence relations for all differential conditions below. The notion of a set of differential and orthogonality conditions (6.2) generalizes corresponding orthogonality conditions assumed in [12, 15]. One could like to understand which part of differential conditions (6.2) are independent. Indeed, in the case with known commutation rules for $\mathcal{C}$-elements and differentials, we use (2.5) to normalize its sequence in the definition (2.6) and in the conditions (6.2) by sending all $d_{\overline{a}}$ to the left with respect to $d_{a}$. Then we use known vanishing rules for powers of differentials. In the case when commutation rules are not known, one might use the independence of powers of differentials. The next result of this paper is the following.
@@
The rules (2.5), the differential conditions (6.2), and a set of maximal orders and powers for all $\mathcal{C}$-elements endow $\mathcal{C}$ with the structure of a multiply graded infinite-dimensional differential algebra with respect to a $\cdot_{l}$-multiplication, $l\geq 0$.
@@
The way of operation with sequences of differential and orthogonality conditions follows from the rules (2.5), Leibniz rule (2.2), as well as from taking intro account maximal orders and powers of differentials and elements, i.e., the assumption that some of $\mathcal{C}$ belong to the ideals with respect to orders $\mathcal{I}(p)$ or powers $\mathcal{I}(q)$. Starting from a particular orthogonality or differential condition, we act consequently by the differentials $d_{a}$ and $d_{\overline{a}}$. Secondly, using the $\mathcal{I}(p)$, $\mathcal{I}(q)$ ideal vanishing properties applied to orthogonality conditions, and by using the completeness property of the complex $\mathcal{C}$ with respect to the $\cdot_{l}$-product, we express particular elements $d^{j_{i}}\phi_{i}$ in terms of other $\mathcal{C}$ elements. With the maximal orders and powers for each particular $\mathcal{C}$-element, a differential or an $\mathcal{C}$-element reaches its maximal order or power, and corresponding summand vanishes. Continuing the process, we finally obtain the full structure of differential conditions. A sequence of relations does not not stop as long as coherence conditions on indices are fulfilled, of until the sequence gives the both side identical zero. In some cases, due to the completeness condition for the $\cdot_{l}$-product, a sequence of differential conditions becomes infinite in a certain branch of the hierarchy. Let us reproduce the general structure of relations following from differential conditions for an expression in the form of a multiple product (2.1). The most general configuration associated to a $\cdot_{l}$-multiple product where several $\mathcal{C}$-elements is with elements that are situated at various places and mixed with the corresponding completions. In our setup, we work with differential conditions containing powers of elements of $\mathcal{C}$ as well as orders of the differentials. The general form of the differential condition is given by (6.2), where $\mathcal{D}_{\bm{J}_{s,i}}\phi_{s,i}$ is of the form (2.6), $1\leq i\leq k\leq l$. By applying the differentials to (6.2) one arrives at further differential and orthogonality relations. Recall, that for each differential condition in this proof, there exist a coherence relation for corresponding $\mathcal{C}$-indices. The sequences of differential and orthogonality conditions derived above together with corresponding coherences relations for indices, taken for all choices of ${\bm{J}}_{s}$, provide the full set of differential and orthogonality relations defining the multiply graded differential algebra. The structure of a resulting differential algebra relations as well as the structure of corresponding closed products given by Theorem 1 depends on which differential or orthogonality condition of a hierarchy we started from. In practice, one can start with a particular set of differential or/and orthogonality conditions. Then the resulting multiple graded differential algebra is a reduction of the full algebra. In addition to that, one can also restrict domain of definitions for gradings for some families of chain-cochain complexes. E.g., one can set $n\geq 0$ instead of $\mathbb{Z}$. This will affect the coherence conditions for corresponding identities. ∎

### v3

@@ 2. Multiple complexes and the main result
Indeed, each a differential identity $(\chi)=0$ from the system (2.8) can be “integrated“ with respect the differentials $d$ or $\overline{d}$ to find a multiple product $(\Gamma)$ such that $d_{a}(\Gamma)=(\chi)=0$. When we say integrated we mean that for an element $\gamma\in\mathcal{C}$ we find an element $\gamma^{\prime}$ such that $\gamma^{\prime}=d_{a}\gamma$. In addition to that relations specified by the identities (2.8) can be used in order to derive closed products which do not follow explicitly from the integration of identities. Let us underline that we work with a multiple complex spaces as grading subspaces of an associative algebra. In particular, the multiple product $\cdot_{l}$ can be introduces for geometric objects of general kinds. Thus, we keep the constructions of this paper independent of commutation relations. In the next three Sections we will provide instructive examples of identities and closed products of the first order of differentials actions. The full description and classification of differential identities and closed products at higher orders and powers of differentials action will be given in a forthcoming paper. The identities and closed products given in Sections (LABEL:porgo)–?? generalize to symmetrized versions with respect to permutations of indices, positions of $\mathcal{C}$-elements, and orders and powers of differentials.
@@ 3. Examples of identities for multiple products
In this Section we illustrate the proof of Theorem ?? by complexity-growing examples. In this and next two Sections we always assume that all completions are closed with respect to corresponding differentials.
@@ 3.1. Two differentials, and the total maximal power two for two kinds of elements
$0$ $=$ $d_{a}\left(\ldots,(\phi)_{i},\ldots,(\psi)_{j},\ldots\right)$
@@ 3.3. Two differentials, identical elements of the total maximal power
For $q(\phi)$ identical elements $\phi$, placed at all $i$-th positions, $1\leq i\leq q(\phi)\in\mathbb{N}$, such that $\phi\in\mathcal{I}(q(\phi))$, i.e., $\phi^{q(\phi)}=0$, equivalently, $(\ldots$, $(\phi)_{1}$, $\ldots$, $(\phi)_{i}$, $\ldots$, $(\phi)_{q(\phi)},\ldots)=0$, (i.e., $\phi$ is placed $q(\phi)$ times in various places in the product). Then,
@@ 3.4. Orders of two differentials of several identical elements
[…]
@@ 3.5. Two differentials, multiple elements
[…]
$0$ $=$ $d_{a}\sum\limits_{i=1}^{k}\left(\ldots,\phi_{1}^{r_{1}},\ldots,\phi_{i}^{q(\phi_{i})},\ldots,(\phi_{k})^{r_{k}},\ldots\right)$ (3.6)
[…]
@@
[…]
@@
One can also take the case of the total maximal power elements. I.e., $\sum_{i=1}^{t}$ $q_{i}$ $(p_{i},\phi_{i})=q$, and $q_{i}(p_{i},\phi_{i})<q(p_{i},\phi_{i})$ i.e., when the sum of powers reaches the critical value, and $\times\left(d_{a}^{p_{i}(\phi_{i})}\phi\right)^{q_{i}(\phi_{i})}\in\mathcal{I}(q)$. Similarly, $\sum_{i=1}^{t}p_{i}(\phi_{i})=p$, and $p_{i}(\phi_{i})<p(\phi_{i})$, i.e., the sum of orders of the differentials is at its maximum, and $\times d_{a}^{p_{i}(\phi_{i})}\phi_{i}\in\mathcal{I}(p)$ Then in both cases $\left(\ldots,\left(d_{a}^{p_{i}(\phi_{i})}\phi_{i}\right)^{q_{i}(p_{i},\phi_{i})},\ldots\right)=0$. When we apply the second differential $d_{\overline{a}}$ to the identity (3.7), all branches of the hierarchy of identities arise with appropriate conditions on orders and powers of differentials.
@@ 4. Examples of a single element closed products
In this Section we provide examples of closed products arising, in particular, from the identities of Section ??. Closed products and their identities are very useful in the theory of completely integrable [2] and exactly solvable [17, 19] systems. Note also possible applications for operads. Note that in the search of products annihilated by a differential, we take only non-vanishing products. To use the annihilation given by maximal powers of $d_{a}\phi_{i}$ in multi-element products after application of a differential $d_{a}$ once, we have to include not just one power of each element but lower powers also. The main idea of a closed product construction is in taking into account changes both in orders of differentials themselves and powers of their action on $\mathcal{C}$-elements. For that reason we have to include not only powers of elements acted by the differentials, but also powers of elements not acted by them. The balance of raising and lowering powers of elements results in extra conditions producing closed products. In some of the examples below we use the transfer of differentials using certain identities from Section ??. By integrating the identities of the form (3.5) one finds sets of advanced closed products.
@@ 4.3. Orders of a single differential, single element, powers of differentials
[…]
@@ 5.4. Powers of orders of multiple differentials, multiple elements
when 1.) there exists $i$, $1\leq i\leq k$, such that $p_{n+1,i}\left(d^{p_{n,i}}_{a_{n,i}}\right$. $\ldots$ $d^{p_{1},i}_{a_{1,i}}$ $\left.\phi\right)$ $\geq$ $p\left(d^{p_{n+1,i}}_{a_{n,i}}\ldots d^{p_{1},i}_{a_{1,i}}\phi\right)$; 2.) if $k\geq 1$, and $({\bf p}_{n+1,i},{\bf a}_{n+1,i})=({\bf p}_{n,j},{\bf a}_{n,j})$, for some $1\leq i\neq j\leq k$.
@@ 6. Multiple-graded differential algebras
where it is clear that $q_{s,i}<q\left(\mathcal{D}_{\bm{J}_{s,i}}\phi_{s,i}\right)$ are less than the maximal powers of the corresponding elements $\mathcal{D}_{\bm{J}_{s,i}}\phi_{s,i}$, and the maximal orders of differentials constituting (as in (LABEL:pirdos)) $\mathcal{D}_{\bm{J}_{s,i}}\phi_{s,i}$ satisfy the recurrence conditions (LABEL:pirdo). We call this a differential condition. Symmetrizing with respect to all choices of $\bm{J}_{s}$} for a fixed $l$, we obtain a set of differential conditions
@@
for $\gamma_{s}\in\mathcal{C}$, $s\in\mathbb{Z}$. With $\mathcal{D}_{\bm{J}_{s}}\gamma_{s}=0$ for some $s$, we call (LABEL:mordo) an orthogonality condition. For each differential condition above of the set (LABEL:mordo) we have the coherence condition for corresponding upper and lower indices of $\mathcal{C}$-spaces, i.e., $\bm{n}$} $\left(\mathcal{D}_{J_{s}}\gamma_{s}\right)$ == $\sum_{i=1}^{l}\left({\bm{n}_{i}}\left(\mathcal{D}_{\bm{J}_{s,i}}\phi_{s,i}\right)-r_{i}\left(\mathcal{D}_{\bm{J}_{s,i}}\phi_{s,i}\right)\right)$, ${\bm{m}}\left(\mathcal{D}_{\bm{J}_{s}}\gamma_{s}\right)=\sum_{i=1}^{l}\left({\bm{m}}_{i}\left(\mathcal{D}_{\bm{J}_{s,i}}\phi_{s,i}\right)\right$. $\left.-{\bm{t}}_{i}\left(\mathcal{D}_{\bm{J}_{s,i}}\phi_{s,i}\right)\right)$. We will skip these coherence relations for all differential conditions below. The notion of a set of differential and orthogonality conditions (LABEL:mordo) generalizes corresponding orthogonality conditions assumed in [13, 16]. One could like to understand which part of differential conditions (LABEL:mordo) are independent. Indeed, in the case with known commutation rules for $\mathcal{C}$-elements and differentials, we use (2.5) to normalize its sequence in the definition (LABEL:pirdos) and in the conditions (LABEL:mordo) by sending all $d_{\overline{a}}$ to the left with respect to $d_{a}$. Then we use known vanishing rules for powers of differentials. In the case when commutation rules are not known, one might use the independence of powers of differentials. The next result of this paper is the following.
@@
The rules (2.5), the differential conditions (LABEL:mordo), and a set of maximal orders and powers for all $\mathcal{C}$-elements endow $\mathcal{C}$ with the structure of a multiply graded infinite-dimensional differential algebra with respect to a $\cdot_{l}$-multiplication, $l\geq 0$.
@@
The way of operation with sequences of differential and orthogonality conditions follows from the rules (2.5), Leibniz rule (LABEL:leibniz), as well as from taking intro account maximal orders and powers of differentials and elements, i.e., the assumption that some of $\mathcal{C}$ belong to the ideals with respect to orders $\mathcal{I}(p)$ or powers $\mathcal{I}(q)$. Starting from a particular orthogonality or differential condition, we act consequently by the differentials $d_{a}$ and $d_{\overline{a}}$. Secondly, using the $\mathcal{I}(p)$, $\mathcal{I}(q)$ ideal vanishing properties applied to orthogonality conditions, and by using the completeness property of the complex $\mathcal{C}$ with respect to the $\cdot_{l}$-product, we express particular elements $d^{j_{i}}\phi_{i}$ in terms of other $\mathcal{C}$ elements. With the maximal orders and powers for each particular $\mathcal{C}$-element, a differential or an $\mathcal{C}$-element reaches its maximal order or power, and corresponding summand vanishes. Continuing the process, we finally obtain the full structure of differential conditions. A sequence of relations does not not stop as long as coherence conditions on indices are fulfilled, of until the sequence gives the both side identical zero. In some cases, due to the completeness condition for the $\cdot_{l}$-product, a sequence of differential conditions becomes infinite in a certain branch of the hierarchy. Let us reproduce the general structure of relations following from differential conditions for an expression in the form of a multiple product (LABEL:tugo). The most general configuration associated to a $\cdot_{l}$-multiple product where several $\mathcal{C}$-elements is with elements that are situated at various places and mixed with the corresponding completions. In our setup, we work with differential conditions containing powers of elements of $\mathcal{C}$ as well as orders of the differentials. The general form of the differential condition is given by (LABEL:mordo), where $\mathcal{D}_{\bm{J}_{s,i}}\phi_{s,i}$ is of the form (LABEL:pirdos), $1\leq i\leq k\leq l$. By applying the differentials to (LABEL:mordo) one arrives at further differential and orthogonality relations. Recall, that for each differential condition in this proof, there exist a coherence relation for corresponding $\mathcal{C}$-indices. The sequences of differential and orthogonality conditions derived above together with corresponding coherences relations for indices, taken for all choices of ${\bm{J}}_{s}$, provide the full set of differential and orthogonality relations defining the multiply graded differential algebra. The structure of a resulting differential algebra relations as well as the structure of corresponding closed products given by Theorem ?? depends on which differential or orthogonality condition of a hierarchy we started from. In practice, one can start with a particular set of differential or/and orthogonality conditions. Then the resulting multiple graded differential algebra is a reduction of the full algebra. In addition to that, one can also restrict domain of definitions for gradings for some families of chain-cochain complexes. E.g., one can set $n\geq 0$ instead of $\mathbb{Z}$. This will affect the coherence conditions for corresponding identities. ∎